OpenStudy (calculusxy):

1. If Lynn can type a page in p minutes, what piece of the page can she do in 5 minutes?

OpenStudy (calculusxy):

@hartnn

hartnn (hartnn):

number of page typed (N) is directly proportional to typing speed(S). so, N =kS 1 page in p minutes , means N=1, S =p so, k = 1/p got this ? so, N=(1/p)S in 5 minutes, means S=5 so, N = 5/p so, in 5 minutes she can do 5/p of the total page . ask if any doubt in any step.